Parth (parthkohli):

Lots and lots of questions. I don't know where to start... there are just so many. Gonna post all of them in this post.

Parth (parthkohli):

Let \(a, b, c\) be integers, and suppose the equation \(f(x) = ax^2 + bx + c = 0\) has an irrational root \(r\). Let \(u = p/q\) be any rational number such that \(|u-r|<1\). Prove that\[\frac{1}{q^2}\le |f(u)| \le K|u-r|\]for some constant \(K\). Deduce that there is a constant \(M\) such that\[\left|r\frac{p}q\right| \ge M/q^2\](This is useful in approximating the non-rational zero of a polynomial by a rational number.)

Parth (parthkohli):

Find the value of the positive integer \(n\) for which the quadratic equation \[\sum_{k=1}^{n} (x+k-1)(x+k) = 10n\]has solutions \(\alpha\) and \(\alpha+1\) for some \(\alpha\).

Parth (parthkohli):

Find all positive integers \(n\) for which the quadratic equation\[a_{n+1}x^2 - 2x \sqrt{\sum_{k=1}^{n+1}a_k^2 }+ \sum_{k=1}^n a_k=0\]has real roots for every choice of real numbers \(a_1, a_2, \cdots, a_{n+1}\)

Parth (parthkohli):

Let the polynomial \(p(z) =z^2 + az + b \) be such that \(a\) and \(b\) are complex numbers and \(|p(z)|=1 \) whenever \(|z| = 1\). Prove that \(a = 0\) and \(b = 0\).

Parth (parthkohli):

Find necessary and sufficient conditions on the coefficients \(a, b, w\) so that the roots of the equations\[z^2 + 2az + b = 0,~ z - w = 0\]are collinear in the complex plane.

Parth (parthkohli):

Let \(ax^2 + bx + c \) be a polynomial with real coefficients such that\[|ax^2+bx+c| \le 1 \]for \(0 \le x \le 1 \). Prove that \(|a| + |b| + |c| \le 17\).

OpenStudy (alekos):

Hi Parth. I'm going to see if I can help you out with these. I've got a couple of days off. How do I save this page so that I can come back to it tomorrow?

Parth (parthkohli):

@alekos Thanks! You can probably bookmark this.

Parth (parthkohli):

Consider the number\[\alpha = \left(\frac{n + \sqrt{n^2 - 4}}{2}\right)^m\]where \(n\ge 2\) and \(m \) are natural numbers. Prove that \[\alpha = \frac{k + \sqrt{k^2 -4 }}{2}\]for some natural number \(k\).

OpenStudy (alekos):

Ok. I'll start in about 3 hours

Parth (parthkohli):

Let \(p(x) = ax^2 + bx + c\) be a polynomial in \(\mathbb R[x]\) such that \(p(\alpha ) \le 1 \) for \(|\alpha | \le 1\). Prove that \(|2a \alpha+b |\le 4 \) for \(|\alpha| < 1 \).

OpenStudy (alekos):

Where are these questions coming from?

Parth (parthkohli):

From a book.

Parth (parthkohli):

I'd happily avoid typing out these questions and rather send you a copy, but a PDF doesn't seem to exist anywhere on the web.

OpenStudy (alekos):

Do you work at a university or something?

Parth (parthkohli):

Let \(p(x)\) be a monic quadratic polynomial over \(\mathbb Z\). Show that for any integer \(n\), there exists an integer \(k\) such that \(p(n) p(n+1) = p(k)\).

Parth (parthkohli):

Nah, just an average high-school student trying to understand how solving olympiad problems works.

OpenStudy (alekos):

Are you thinking of entering?

Parth (parthkohli):

I started out too late, so the best I'll do by learning this would be personal satisfaction. Nothing else.

Parth (parthkohli):

Prove the division algorithm in \(\mathbb R[x]\): given polynomials \(a(x)\) and \(b(x)\) in \(\mathbb R[x]\), there are unique polynomials \(q(x)\) and \(r(x)\) such that \(a(x) = b(x)q(x) + r(x)\) where either \(r(x)\) is the zero polynomial or \(\deg r(x) < \deg b(x)\).

OpenStudy (mom.):

for the 4th one- \[a=x_{1}+ib_{1}\]\[b=x_{2}+ib_{2}\] |z| can be 1 so z can take these values- 1,-1, i we have 3 cases- 1)z=1 \[1=|1+x_{1}+ib_{1}+x_{2}+ib_{2}|\]\[1=|(1+x_{1}+x_{2})+(y_{1}+y_{2})i|\]\[1=\sqrt{(1+x_{1}+x_{2})^2+(y_{1}+y_{2})^2}\] from here we can easily conclude that x1,x2,y1,y2 are 0 and hence a,b are 0 2)z=-1 its similar and we will get this thing at last- \[1=\sqrt{(1-x_{1}+x_{2})^2+(y_{2}-y_{1})^2}\] here also u can get the same conclusion 3)z=i \[1=|1+(x_{1}+iy_{1})i+(x_{2}+y_{2}i)|\] \[1=\sqrt{(y_{2}+x_{1})^2 +(1+x_{2}-y_{1})^2}\] here also u can conclude the same

OpenStudy (mom.):

oops there will be one more case of z= -i u can follow the same conclusion there :)

Parth (parthkohli):

There are actually infinitely many cases! I guess we'll have to work for the general case...

OpenStudy (baru):

yep, infinite cases, take any \(\theta\) on the unit circle of complex plane

Parth (parthkohli):

But still, that was the correct initiation.

OpenStudy (mom.):

that jst slipped outta my mind sry yeah we must consider this thing- \[Z=\sqrt{(x_{1}^2+y_{1}^2)}e^{i \theta}\] so \[1=|1+(\sqrt{(x_{1}^2+y_{1}^2)}e^{i \theta})i +\sqrt{(x_{2}^2+y_{2}^2)}e^{i \theta}|\]\[1=|1+(\sqrt{(x_{1}^2+y_{1}^2)}e^{i \theta})i +\sqrt{(x_{2}^2+y_{2}^2)}e^{i \theta}|\] now what:/

ganeshie8 (ganeshie8):

Find the value of the positive integer \(n\) for which the quadratic equation \[\sum_{k=1}^{n} (x+k-1)(x+k) = 10n\]has solutions \(\alpha\) and \(\alpha+1\) for some \(\alpha\). Solution : Since \(\alpha\) and \(\alpha+1\) are the roots, \[\sum_{k=1}^{n} (\alpha +k-1)(\alpha +k) = 10n\tag{1}\] \[\sum_{k=1}^{n} (\alpha+k)(\alpha+k+1) = 10n\tag{2}\] \((2)-(1) \implies\) \[\sum_{k=1}^{n} (\alpha+k) = 0\tag{3}\] \((2)+(1) \implies\) \[\sum_{k=1}^{n} (\alpha+k)^2 = 10n\tag{4}\] solving \((3)\) and \((4)\) gives \(n=11\)

Parth (parthkohli):

Jesus, I didn't even try solving that problem. Thank you so much!

ganeshie8 (ganeshie8):

These problems are really nice... each seem to have a clever solution..

Parth (parthkohli):

Wait, have you solved all of them?

ganeshie8 (ganeshie8):

nope, still working on #1

OpenStudy (ivysleague):

Could I have a hand with geometry?

OpenStudy (readergirl12):

Um..Just looking at those "equations" gave me a headache.. o.e

OpenStudy (anonymous):

umm your very worderous how much do you wanna know your brain is gonna be sooo big dude

Parth (parthkohli):

I'm gonna give a go at one of those problems.\[z = w\tag {1}\]For \(z^2 +2az+b = 0 \) to have roots that are collinear with those of the first equation, clearly the roots are scalar multiples of \(w\). Let \(\lambda w, -2a - \lambda w\) be the roots. We have to eliminate \(\lambda\). How?!

Parth (parthkohli):

Or not. Collinearity does not imply that three complex numbers are scalar multiples of each other.

Parth (parthkohli):

If \(z_1, z_2\) are the roots, then\[w - z_1 = \lambda(z_1 - z_2)\]\[z_1 + z_2 = -2a\]\[z_1z_2=b\]Eliminating \(z_1, z_2, \lambda\) is what we have to do. OK.

OpenStudy (empty):

I wanna TL;DR this lol which ones are unanswered still?

Parth (parthkohli):

All except the one ganeshie answered. Also you need the discriminant in this question: Find all positive integers \(n\) for which the quadratic equation\[a_{n+1}x^2 - 2x \sqrt{\sum_{k=1}^{n+1}a_k^2 }+ \sum_{k=1}^n a_k=0\]has real roots for every choice of real numbers \(a_1, a_2, \cdots, a_{n+1}\)

Parth (parthkohli):

\[\sum_{k=1}^{n+1}a_k^2 \ge a_{n+1}\sum_{k=1}^n a_k\]Some kinda inequality? I don't know.

Parth (parthkohli):

exactly... but how do I apply that?

OpenStudy (empty):

I've never really understood this, but it's supposedly pretty important https://en.wikipedia.org/wiki/Polynomial_remainder_theorem#Proof

Parth (parthkohli):

yes, but proving the division algorithm itself seems like a huge deal.

ganeshie8 (ganeshie8):

http://www.artofproblemsolving.com/wiki/index.php/Power_Mean_Inequality gives \[\dfrac{1}{n}\sum_{k=1}^{n}a_k^2 \ge \left(\dfrac{1}{n}\sum_{k=1}^n a_k\right)^2\]

Parth (parthkohli):

thank you sire

Parth (parthkohli):

hmm, the sum indices are a little different. how do I handle all that?

Parth (parthkohli):

@ganeshie8 could you post it again

ganeshie8 (ganeshie8):

I think we need to massage it a bit more before applying the inequality..